This question tests your knowleige of encryption and decryption wsing the RSA method. The numbers insole ed are delilerotely chosen to be mall enough to focus on gown understanding without eroessive colculatioms. Alice and Bob decide to use an RSA eryptosystem with public key (77,13) for communication. (a) Alice wants to send Bob the message m 22. Determine Alices ciphertext c. (b) Determine Bobs private (c) Carry out Bobs decryption of Alies ciplaertext and conupare his result with Alices message
Answer #1

1.Lets assume the two keys of alice and bob be p and q respectively

P=77, q=13

Now according to RSA algorithm

N=p*q = 77*13 = 1001

φ = (p-1)(q-1) = 76*12 = 912

Now we choose an integer e such that 1 < e < φ(n) and e and φ (n) are coprime. Let e =11

Compute a value for d such that (d * e) % φ(n) = 1. One solution is d = 83

[( 83* 11) % 912 = 1]

Public key is (e, n) => (11,1001)

2.Private key is (d, n) => (83, 1001)

3.Compare of results for bob and alice

The encryption of m = 29 is c = 29^11 % 1001 = 679

The decryption of c = 679 is m = 679^77 % 1001 = 29

Please note it is possible solution for RSA algorithm .However there can be multiple solution existing